Hence, write down the equation with y as the subject (in terms of x). The value of m and q must be indicated


1
Expert's answer
2022-03-07T21:22:01-0500

Solution


Consider the equation,


"4x^ny - 2y -x^m=0"


Here, "m" and "n" are constants.


To make the "y" the subject and write in terms of ""x"", we proceed as follows.


"4x^ny - 2y -x^m=0"


"4x^ny - 2y=x^m"


"y(4x^n - 2) =x^m"


"y =\frac{x^m}{4x^n - 2}"


"y(x) =\frac{x^m}{4x^n - 2}"


Hence ""y"", in terms of ""x" " is


"{\color{Blue} y(x) =\frac{x^m}{4x^n - 2}\\}"